OpenStudy (anonymous):

A racing car travels at an average speed of 90km/h around a 12 km track. (a) how long will it take to make 12 round of the track?

OpenStudy (anonymous):

V = s/t so t = s/v = 12/ 90 for 1 round Finally for 12 round t = 12 x 12/90 = 1.6 Hour

OpenStudy (anonymous):

What is V?

OpenStudy (anonymous):

Could you re read what your typed. i don't understand.

OpenStudy (anonymous):

v is speed or velocity...,

OpenStudy (anonymous):

V = s/t I'm looking for the " Time" not " Speed"

OpenStudy (anonymous):

v = s/t where v is speed or velocity, s is distance or displacement, and t is time so t = s/v = 12/ 90 for 1 round Finally for 12 round t = 12 x 12/90 = 1.6 Hour

OpenStudy (anonymous):

yes i'm looking for time too.., look carefully t = s/v = 12/ 90 for 1 round (where t is time) Finally for 12 round t = 12 x 12/90 = 1.6 Hour (where t is time)

OpenStudy (anonymous):

t = s/v = 12/ 90 for 1 round T= Time? S= Speed V= Velocity?

OpenStudy (anonymous):

yes of course ur right! :)

OpenStudy (anonymous):

Of yes i get it now

OpenStudy (anonymous):

congrats ;)